OpenStudy (anonymous):

Rule of derivatives: y = 3x + 7 √x

OpenStudy (campbell_st):

rewrite using indices \[y = 3x + 7x^{\frac{1}{2}}\] now you can differentiate

OpenStudy (anonymous):

Where does the 1/2 come from?

OpenStudy (campbell_st):

well the square root... the index law is \[\sqrt{x} = x^{\frac{1}{2}}\] or in general the nth root is \[\sqrt[n]{x} = x^{\frac{1}{n}}\] where n is an positive integer